OpenStudy (anonymous):

how do you write 3.04 as a mixed numer and reduce?

hero (hero):

Hint: \[3\frac{4}{100}\]

OpenStudy (wwe123):

3.04 = \[3\frac{ 4 }{ 100 }\] = \[3\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

you guys really really explian better than ma teacher

OpenStudy (wwe123):

sorry \[3\frac{ 1 }{ 25}\]

OpenStudy (anonymous):

is like 400 becomes a quarter or 25 sorry for my logic?

OpenStudy (unklerhaukus):

\[3\frac4{100}=3\frac{4}{4\times25}=3\frac{\cancel 4}{\cancel 4\times25}=3\frac 1{25}\]